For the following problem, the slope and one point on a line are given. Use the point-slope form to find the equation of the line. (Write your answers in slope-intercept form.)

(-8,-1), m = -1/4

Found 2 solutions by rfer, mananth:
Answer by rfer(16322)
Answer by mananth(16946)

the equation is given by
(y-y1)=m(x-x1)
..
(-8,-1), m = -1/4
..
y-(-1)=(-1/4)*(x-(-8))
y+1=(-1/4)(x+8)
y+1=-x/4 -2
y=(-x/4)-3
...
m.ananth@hotmail.ca